# Do We Run Large-scale Multi-Robot Systems on the Edge? More Evidence for Two-Phase Performance in System Size Scaling

### v1

@@ I Introduction
Swarm robotics achieves scalability by enforcing a decentralized approach based on self-organizing behaviors, local robot-robot interactions, and simple neighbor-based communication [4]. Several studies have measured a certain intuitive swarm performance (e.g., number of collected items during a defined time interval in a foraging task [6]), usually over swarm density $\rho=N/A$, for swarm size $N$ and a bounded operation area of size $A$ [7, 8, 9, 3]. If area size $A$ is fixed, scaling swarm size $N$ effectively modulates the swarm density. Almost all studies that measure mean swarm performance over swarm density (e.g., see [10]) show similar curves characterized by three stages: an initial increase with system size, reaching peak performance (at a critical system size $N_{c}$), and for $N>N_{c}$ decreasing system performance. Interestingly, a diversity of systems from different fields of research show similar scalability results (e.g., see [11, 12, 13]). While measuring mean swarm performance is a good approach for an initial performance analysis, it is likely to hide an important property. The measured swarm performance may be bimodal showing two phases: an upper phase with high performance (i.e., each robot operates efficiently most of the time), a lower phase with minimal performance (i.e., most robots trapped in congestion etc.), and almost nothing in between [14]. This is, however, a potentially dangerous condition as swarm robotics systems often show linear or even super-linear performance increases for low swarm densities. By optimizing for performance, one may unknowingly put the system on the edge, where model predictions and real-world experience may catastrophically diverge. In addition, the robotic system is likely to have a long transient at the critical system size which may have two implications. First, the system can actually be run on ‘borrowed time’ if the transient is longer than what is relevant in applications (i.e., potentially contradicting mathematical predictions built on assumed convergence). Second, the system at runtime initially shows good performance but then suddenly breaks down completely.
@@ II-B Object Clustering
In the object clustering task a swarm of robots manipulate objects, either bringing them to a pre-defined goal position, or aggregating them in a self-organized manner (see Fig. 0(b)). If the goal is pre-defined, then this task becomes synonymous with foraging [17]. We test the lasso method for object clustering as its performance as a function of swarm size has already been partially explored [18]. This method relies on an external scalar field which acts to guide the robots around the environment while nudging objects towards the global minimum of the field, located at the goal.
The control algorithm operates independently on each robot. The most outlying visible puck is identified and the value of this puck’s position on the scalar field is determined. A control law drives the robot to reach and follow the contour line defined by this value. We observe that the robots tend to occupy the same contour line and travel clockwise around it, nudging pucks inwards toward the goal. Thus, a ‘lasso’ is formed that tightens around the objects until they are forced to the goal position. Performance was found to increase with swarm size up to a point, then decrease due to spatial interference [18]. The results depicted in Fig. 0(e) were collected in a custom simulator which models two-dimensional physics111open-source code repository: https://github.com/avardy/cwaggle_lasso, video: https://youtu.be/_KOU5SzpQBg..
@@ III-A Warehouse
At low arrival rates, the system does not suffer from congestion and the throughput increases linearly with the arrival rate, as expected for a system with low density. At an arrival rate of about $10^{4}$ units per hour interference effects lead to some experiments in which the throughput drops considerably. This drop of throughput is caused by a congestion in the warehouse, when too many AFLEs interfere with each other. When two AFLEs are trying to traverse the same rail segment, they need to negotiate the traversal order and AFLEs with lower priority wait until AFLEs with higher priorities are passed. With increased arrival rate, the number of experiments with congestion increases until all repetitions suffer from congestion. It is important to note that there is a range of arrival rate values for which part of the experiments operated properly, with maximum performance, and the remaining experiments collapsed into complete congestion with performance close to zero. As a result, the distribution of the throughput splits into two phases, one high performing and one low performing.
@@ III-B Object Clustering
Figure 2: The performance plot from Fig. 0(e) augmented with snapshots of the final arena configuration. The robots are displayed in red. Pucks are shown in green. The goal position is depicted at top-left.
Fig. 0(e) shows the performance of simulated swarms performing the object clustering task for increasing swarm sizes, from one to twenty robots. For each swarm size, we evaluated the performance $1000$ times. Initially, all robots and objects were randomly distributed in the arena. The performance is the normalized average distance of all objects to the target location. A performance measure of 00 is the ideal, but is not attainable due to physical interference between objects.
@@ IV-A Queues
A simple approach of modeling the two-phase situation of swarm performance can be based on queueing theory. An intuitive and sophisticated model could be based on queueing networks to explicitly represent the shared resources, such as space and radio bandwidth (possibly also truncated queues). However, a simpler approach is possible by an unorthodox change of the standard M/M/1 queue with exponential interarrival-time distribution $\lambda\exp{(-\lambda t)}$, exponential service time $\mu\exp{(-\mu t)}$, one channel, infinite capacity. Key change to the standard model is that we make the service time dependent on the current queue length $N_{q}(t)$ (i.e., a cost on keeping the queue) by setting mean service time $1/\mu(t)=1/(c_{1}\log(c_{2}N_{q}(t))+\mu_{c})$ for arbitrary constants $c_{1}=0.3$ and $c_{2}=0.2$. For the following example we set $\mu_{c}=1$. We simulate the queue for mean arrival rates $\lambda\in[0.4,0.9]$ (500 samples each) for $2\times 10^{6}$ time steps and exclude data from the first $2\times 10^{6}$ time steps as transient. Resulting mean throughput and mean queue length are shown in Fig. 2(a) and a histogram of the throughput in Fig. 2(b). It is to note that the curves shown in Fig. 2(a) resemble those predicted by the population model (see Fig. 5a). This queuing model is able to represent the linear increase and the breakdown of the system at a critical arrival rate $\mu_{c}\approx 0.65$.
@@ IV-B Population model and nonlinear dynamics analysis
In previous work, we have presented a population model that grasps most of the relevant system features concerning scalability [3]. In this alternative approach to model scalability we assume that each robot can be assigned to one of three abstract robot states relevant for scalability analysis. The system dynamics is based on $N$ robots that transit between these three operational states. We assume robots can be in state solo (operating alone without sharing communication and/or space resources with others), in state grupo (sharing resources with other robots), or in state fermo (interfering with others in the use of resources). The model describes through seven transition rates ($k_{1},\dots,k_{7}$) the probability that robots change their state as a consequence of independent actions or interactions with other robots. As in any population model we make assumptions, for example, that it is more likely for a solo robot to approach a solo robot if the solo state currently has a high percentage in the swarm. We assume also that two solo robots approaching each other trigger transitions to grupo robots and similar for any other interaction of two robots (see Fig. 4).
@@
Figure 5: Stationary states of Eq. (1) under variation of $N$ for (a) $k_{1}=0.005$ and (b) $k=0.001$, other parameters as in Eq. (2). (b) Domain between the saddle-node bifurcation points $N^{(1)}_{\text{\tiny SN}}$, $N^{(2)}_{\text{\tiny SN}}$ associated with bi-stability, the non-physical domain outside of $0\leq s,f,g\leq N$ is shown in gray; (c) transient time $T_{\text{trans}}$ for fixed initial state $(s_{0},f_{0})=(0,0)$ and convergence to stationary state $(s_{1}^{*},f_{1}^{*})$ with accuracy $10^{-12}$; $k_{1}=0.005$. Insets show the indicated rectangles magnified.
@@
For a quick understanding of this model, we have analyzed the stationary system behavior. Fig. 5a shows the stationary states $(s^{*},f^{*})$ of system (1) at the parameter set specified in Eq. (2) under variation of swarm size $N$. For completeness, the resulting value $g^{*}=N-s^{*}-f^{*}$ due to robot conservation is shown, too. Fig. 5a shows, how for small swarm sizes $N$, the number of solo robots $s^{*}$ grows almost linearly (most robots remain in solo, few in fermo and grupo). For increasing swarm size $N$, a sudden change of the system’s configuration occurs. The number of solo robots reaches a maximum (here, $N=31$) and then decrease rapidly. Simultaneously, we get more fermo robots, resulting eventually in their predominance. From the perspective of nonlinear dynamics, the shifted location of the asymptotically stable fixed point is not a bifurcation as the topological structure of the state space remains unchanged under parameter variation. However, in any practical application the sudden decrease of solo robots would have a significant impact on system performance.
@@ V Discussion and Conclusion
We have presented three multi-robot systems that show the characteristic two-phase performance signature where, for a critical swarm size, the phases of high and low performance coexist. We have presented that for simulations of the AFLE robot system by EMHS GmbH. Hence, our findings clearly show that not only real-world implementations of robot swarms but also real-world applications of actual productive systems in existing industries show this same property in their system size scaling (or in this case scaling with arrival rates). We have also presented simulation results of an object clustering scenario that was previously implemented on real robots, showing the validity of the simulation [18]. Also in this more academic example from swarm robotics, we observe the same two-phase performance signature. As a last example, we have reminded the reader of simulation results for the well-known emergent taxis scenario that were published before [14]. This original finding inspired the presented research. To point to a possible interpretation of these results, we have presented two modeling options based on queueing theory and population models. The adapted M/M/1 queue is able to catch most of the discussed scalability features despite its simplicity. Our key contribution in terms of modeling, however, is the population model and especially its interpretation based on nonlinear dynamics. These interpretations point to three important findings and issues. (a) There is indeed a bifurcation in the system and rapid changes in performance with small changes in swarm size should be expected. (b) In the vicinity of bifurcation points we should indeed expect extremely increased transient times. (c) We should expect hysteresis effects depending on swarm size that is increased or decreased online while operating. Swarm robotics systems that are optimized for swarm performance may unknowingly be put on the edge with the effect of unexpected, catastrophic performance breakdowns. Long transients may either obscure this effect or may buy the robot swarm time that is possibly long enough to survive in high-performing states for the whole mission time. Future research will hopefully provide more evidence of whether the two-phase performance signature in scalability of swarm robotics is real and relevant even in industrial applications.

### v2

@@ I Introduction
Swarm robotics achieves scalability by enforcing a decentralized approach based on self-organizing behaviors, local robot-robot interactions, and simple neighbor-based communication [4]. Several studies have measured a certain intuitive swarm performance (e.g., number of collected items during a defined time interval in a foraging task [6]), usually over swarm density $\rho=N/A$, for swarm size $N$ and a bounded operation area of size $A$ [7, 8, 9, 3]. If area size $A$ is fixed, scaling swarm size $N$ effectively modulates the swarm density. Almost all studies that measure mean swarm performance over swarm density (e.g., see [10]) show similar curves characterized by three stages: an initial increase with system size, reaching peak performance (at a critical system size $N_{c}$), and for $N>N_{c}$ decreasing system performance. Interestingly, a diversity of systems from different fields of research show similar scalability results (e.g., see [11, 12, 13]). While measuring mean swarm performance is a good approach for an initial performance analysis, it is likely to hide an important property. The measured swarm performance may be bimodal showing two phases: an upper phase with high performance (i.e., each robot operates efficiently most of the time), a lower phase with minimal performance (i.e., most robots trapped in congestion etc.), and almost nothing in between [14]. This is, however, a potentially dangerous condition as swarm robotics systems often show linear or even super-linear performance increases for low swarm densities. By optimizing for performance, one may unknowingly put the system on the edge, where model predictions and real-world experience may catastrophically diverge. Also, the robot system likely has a long transient at critical system size with two implications. First, the system can actually be run on ‘borrowed time’ if the transient is longer than what is relevant in applications (i.e., potentially contradicting mathematical predictions built on assumed convergence). Second, the system at runtime initially shows good performance but then suddenly breaks down completely.
@@ II-B Object Clustering
In the object clustering task a swarm of robots manipulate randomly placed objects, either bringing them to a pre-defined goal position, or aggregating them in a self-organized manner (see Fig. 1(b)). If the goal is pre-defined, then this task becomes synonymous with foraging [17]. We test the lasso method for object clustering as its performance as a function of swarm size has already been partially explored [18]. This method relies on an external scalar field which acts to guide the robots around the environment while nudging objects towards the global minimum of the field, located at the goal.
The control algorithm operates independently on each robot. The most outlying visible puck is identified and the value of this puck’s position on the scalar field is determined. A control law drives the robot to reach and follow the contour line defined by this value. We observe that the robots tend to occupy the same contour line and travel clockwise around it, nudging pucks inwards toward the goal. Thus, a ‘lasso’ is formed that tightens around the objects until they are forced to the goal position. Performance was found to increase with swarm size up to a point, then decrease due to spatial interference [18]. Fig. 1(e) shows results collected in a custom simulator which models two-dimensional physics.111open-source code repository: https://github.com/avardy/cwaggle_lasso, video: https://youtu.be/_KOU5SzpQBg.
@@ III-A Warehouse
At low arrival rates, the system does not suffer from congestion and the throughput increases linearly with the arrival rate, as expected for a system with low density. At an arrival rate of about $10^{4}$ units per hour, interference effects may cause the throughput to drop considerably in few runs. This is caused by a congestion in the warehouse, when too many AFLEs interfere with each other. When two AFLEs are trying to traverse the same rail segment, they need to negotiate the traversal order and AFLEs with lower priority wait until AFLEs with higher priorities are passed. With increased arrival rate, the number of experiments with congestion increases until all repetitions suffer from congestion. Note that there is a range of arrival rate values for which part of the experiments operated properly, with maximum performance, and the remaining experiments collapsed into complete congestion with performance close to zero. As a result, the distribution of the throughput splits into a high-performing and a low-performing phase.
@@ III-B Object Clustering
Figure 2: The performance plot from Fig. 1(e) augmented with snapshots of the final arena configuration. Robots are red. Pucks are green. The goal position is depicted at top-left.
Fig. 1(e) shows the performance of simulated swarms performing the object clustering task, for swarm sizes from one to twenty robots. For each swarm size, we evaluated the performance $1000$ times. The performance is the normalized average distance of all objects to the target location. A performance measure of 00 is the ideal, but is not attainable due to physical interference between objects.
@@ IV-A Queues
A simple approach of modeling the two-phase situation of swarm performance can be based on queueing theory. An intuitive and sophisticated model could be based on queueing networks to explicitly represent the shared resources, such as space and radio bandwidth (possibly also truncated queues). However, a simpler approach is possible by an unorthodox change of the standard M/M/1 queue with exponential interarrival-time distribution $\lambda\exp{(-\lambda t)}$, exponential service time $\mu\exp{(-\mu t)}$, one channel, infinite capacity. Key change to the standard model is that we make the service time dependent on the current queue length $N_{q}(t)$ (i.e., a cost on keeping the queue) by setting mean service time $1/\mu(t)=1/(c_{1}\log(c_{2}N_{q}(t))+\mu_{c})$ for arbitrary constants $c_{1}=0.3$ and $c_{2}=0.2$. For the following example we set $\mu_{c}=1$. We simulate the queue for mean arrival rates $\lambda\in[0.4,0.9]$ (500 samples each) for $2\times 10^{6}$ time steps and exclude data from the first $1\times 10^{6}$ time steps as transient. Resulting mean throughput and mean queue length are shown in Fig. 3(a) and a histogram of the throughput in Fig. 3(b). It is to note that the curves shown in Fig. 3(a) resemble those predicted by the population model (see Fig. 5a). This queuing model is able to represent the linear increase and the breakdown of the system at a critical arrival rate $\mu_{c}\approx 0.65$.
@@ IV-B Population model and nonlinear dynamics analysis
In previous work, we have presented a population model that grasps most of the relevant system features concerning scalability [3]. Key advantages are that it is a general, unified model that abstracts and describes scalability patterns across different systems, including parallel supercomputers, robot swarms, and wireless sensor networks, based on the microscopic interactions between units. In this alternative approach to model scalability we assume that each robot can be assigned to one of three abstract robot states relevant for scalability analysis. The system dynamics is based on $N$ robots that transit between these three operational states. We assume robots can be in state solo (operating alone without sharing communication and/or space resources with others), in state grupo (sharing resources with other robots), or in state fermo (interfering with others in the use of resources). The model describes through seven transition rates ($k_{1},\dots,k_{7}$) the probability that robots change their state as a consequence of independent actions or interactions with other robots. As in any population model we make assumptions, for example, that it is more likely for a solo robot to approach a solo robot if the solo state currently has a high percentage in the swarm. We assume also that two solo robots approaching each other trigger transitions to grupo robots and similar for any other interaction of two robots (see Fig. 4).
@@
Figure 5: Stationary states of Eq. (1) under variation of $N$ for (a) $k_{1}=0.005$ and (b) $k_{1}=0.001$, other parameters as in Eq. (2). (b) Domain between the saddle-node bifurcation points $N^{(1)}_{\text{\tiny SN}}$, $N^{(2)}_{\text{\tiny SN}}$ associated with bi-stability, the non-physical domain outside of $0\leq s,f,g\leq N$ is shown in gray; (c) transient time $T_{\text{trans}}$ for fixed initial state $(s_{0},f_{0})=(0,0)$ and convergence to stationary state $(s_{1}^{*},f_{1}^{*})$ with accuracy $10^{-12}$; $k_{1}=0.005$. Insets show the indicated rectangles magnified.
@@
For a quick understanding of this model, we have analyzed the stationary system behavior. Fig. 5a shows the stationary states $(s^{*},f^{*})$ of system (1) at the parameter set specified in Eq. (2) under variation of swarm size $N$. For completeness, the resulting value $g^{*}=N-s^{*}-f^{*}$ due to robot conservation is shown, too. Fig. 5a shows, how for small swarm sizes $N$, the number of solo robots $s^{*}$ grows almost linearly (most robots remain in solo, few in fermo and grupo). For increasing swarm size $N$, a sudden change of the system’s configuration occurs. The number of solo robots reaches a maximum (here, $N=31$) and then decrease rapidly. Simultaneously, we get more fermo robots, resulting eventually in their predominance. From a nonlinear dynamics perspective, the shifted location of the asymptotically stable fixed point is not a bifurcation as the topological structure of the state space remains unchanged under parameter variation. However, in practical applications the sudden decrease of solo robots would have a significant impact on system performance.
@@ V Discussion and Conclusion
We have presented three multi-robot systems that show the characteristic two-phase performance signature where, for a critical swarm size, the phases of high and low performance coexist. We have presented simulations of the AFLE robot system by EMHS GmbH, showing that not only robot swarms but also real-world applications of actual productive systems in existing industries show this same property in their system size scaling. We have also presented simulation results of an object clustering scenario that was previously implemented on real robots, showing the validity of the simulation [18]. Also in this more academic example, we observe the same two-phase performance signature. As a last example, we have reminded the reader of simulation results for the well-known emergent taxis scenario that were published before [14]. This original finding inspired the presented research. To point to a possible interpretation of these results, we have presented two modeling options based on queueing theory and population models. The adapted M/M/1 queue is able to catch most of the discussed scalability features despite its simplicity. Our key contribution in terms of modeling, however, is the population model and especially its interpretation based on nonlinear dynamics. These interpretations point to three important findings and issues. (a) There is indeed a bifurcation in the system and rapid changes in performance with small changes in swarm size should be expected. (b) In the vicinity of bifurcation points we should indeed expect extremely increased transient times. (c) We should expect hysteresis effects depending on swarm size that is increased or decreased online while operating. Swarm robotics systems that are optimized for performance may unknowingly be put on the edge with the effect of unexpected, catastrophic performance breakdowns. Long transients may either obscure this effect or may buy the robot swarm time that is possibly long enough to survive in high-performing states for the whole mission time. Future research will hopefully provide more evidence of whether the two-phase performance signature in scalability of swarm robotics is real and relevant even in industrial applications.
Acknowledgment
HH thanks Pawel Romanczuk for the valuable discussions.